I can't seem to understand this problem..

Please help me, I'm beging someone ....PLEASE

F(X)=X^2-2, THEN, F(X+2)-F(X)=????

Answer by ionesco(12)

f(x)=x^2-2 then f(x+2)=(x+2)^2-2 then
f(x+2)-f(x)= (x+2)^2-2-(x^2-2)=x^2+4x+4-2-(x^2-2)=4x-6